the sum of the squarees of two consecutive positive odd intergers is one hundred thirty. Find the two intergers.
smaller interger=
larger interger=

Answer by stanbon(75887)

the sum of the squares of two consecutive positive odd integers is one hundred thirty. Find the two integers.
-------
1st: 2x-1
2nd: 2x+1
----
Equation:
(2x-1)^2 + (2x1)^2 = 130
4x^2 -4x + 1 + 4x^2 + 4x + 1 = 130
----
8x^2 + 2 = 130
8x^2 = 128
x^2 = 16
Positive solution:
x = +4
1st: 2x-1 = 7
2nd: 2x+1 = 9
====================
Cheers,
Stan H.
===============